The car wheel makes 180 rpm. Wheel radius 30cm. What is the speed of the car?


1
Expert's answer
2021-03-08T08:30:32-0500

Let's first find the distance covered in one revolution:


"d=2\pi r=2\pi\cdot30\ cm=188.4\ cm."

Finally, we can find the velocity of the car:


"v=180\ \dfrac{rev}{min}\cdot188.4\ \dfrac{cm}{rev}\cdot\dfrac{60\ min}{1\ h}\cdot\dfrac{1\ km}{10^5\ cm}=20.35\ \dfrac{km}{h}."